ROTATIONAL SPECTRUM OF PROPARGYL ALCOHOL DIMER

DEVENDRA MANI and \underline{E. ARUNAN}, Department of Inorganic and Physical Chemistry, Indian Institute of Science Bangalore, India-560012.

Propargyl alcohol is a molecule of interest to astrophysics as well as combustion studies. Rotational-tunneling spectra of propargyl alcohol monomer is well known and shows that the molecule exists in gauche form^{a}^{b}. Recently we reported microwave spectra of Ar...propargyl alcohol complex^{c}. Propargyl alcochol exists in gauche form in the complex as well. In this study we have recorded pure rotational spectra of propargyl alcohol dimer between 4-13 GHz range.A total of 47 transitions, 24 a-type, 16 b-type and 7 c-type, have been observed and fitted with semi rigid rotor asymmetric top hamiltonian. The fitted rotational constants are: A = 2321.83323(47) MHz, B = 1150.47726(24) MHz and C = 1124.89000(20) MHz. The standard deviation for the fit is 2.5 kHz. The experimental rotational constants are very close to the structure predicted by ab-initio calculations in which two gauche-propargyl alcohol moieties are in three point contact stabilized by O-H...O, O-H...pi and C-H...pi interactions. Few transitions for duterated isotopologues of the dimer have also been observed and search for the remaining transitions is in progress. Details will be presented in the talk.

aE. Hirota,J. Mol. Spectrosc. 26 (1968) 335-350.

bJ.C. Pearson, B.J. Drouin, J. Mol. Spectrosc. 234 (2005) 149-156.

cD. Mani, E. Arunan, ChemPhysChem 14 (2013) 754-763.